half life period of a first order chemical is 6.93 minuts.the time require for the completion of 99% of this chemical reaction will be..

Grade:11

1 Answers

Sanyu daver
6 years ago
Half life of first order reaction is = 0.693/kTherefore k= 0.693/6.93 = 0.1K=0.1- completion of 99% of reaction means that the final value of reactant has reduced to 1/100 ((1-99/100 =1/100)of its initial valueAs for first order reaction ;K=2.303/t ×log(initial value of reactant/final value of reactant)K=2.303/t×log100T= 2.303/0.1 × 2log 10 = 2.303×2/0.1 =4.606/0.1 = 46.06minutes